OpenStudy (anonymous):

A bag contains 4 white marbles, 3 black marbles, and 5 red marbles. Find the probability of picking a red marble

OpenStudy (diyadiya):

There is a total of 4 + 3 + 5 =12 marbles number of red marbles = 5 So probability of picking Red Marble = 5/12

OpenStudy (diyadiya):

did you understand? @4everjames

OpenStudy (anonymous):

@Diyadiya yes!!!! ur awesome!!! but this one.is hard because u have to pick another one..A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a black marble, replacing it, and then picking a black marble?

OpenStudy (diyadiya):

so tell me the total number of marbles here

OpenStudy (anonymous):

15

OpenStudy (diyadiya):

Yes so what is the probability of picking a black marble?

OpenStudy (diyadiya):

can you tell me the total number of black marbles?

OpenStudy (anonymous):

ohhhh 4/15

OpenStudy (anonymous):

omg right because u replace it! meaning u have the same chance to get it!!

OpenStudy (apoorvk):

These are the statements of the same question or is this one a different? @4everjames

OpenStudy (apoorvk):

*one's No, james, doesn't work that way actually. one moment

OpenStudy (diyadiya):

Yes!! so we pick it and replace it back now again when you pick for the next time the probabability remains 4/15 and to find the probability of these two events together ( probability of picking a black marble, replacing it, and then picking a black marble ) multiply it .

OpenStudy (apoorvk):

No worries, we'll get you back on the roll soon! And, like our word-filter? :D

OpenStudy (diyadiya):

multiply it means multiply both probabilities

OpenStudy (apoorvk):

I hope this is not combined with the 'bag' posted at the top.

OpenStudy (diyadiya):

lol

OpenStudy (anonymous):

@apoorvk hahaahha the word filter lol

OpenStudy (diyadiya):

@4everjames Answerrrrrr???

OpenStudy (anonymous):

4/15

OpenStudy (diyadiya):

well read my last part again "" to find the probability of these two events together ( probability of picking a black marble, replacing it, and then picking a black marble ) multiply it(multiply it means multiply both probabilities) .""

OpenStudy (anonymous):

8/75?

OpenStudy (diyadiya):

how ? ;)

OpenStudy (anonymous):

multiplied both probabilities

OpenStudy (diyadiya):

what are the two probabilities here?

OpenStudy (apoorvk):

4 times 4 on earth was 16, and 15 times 15 was two two five!, so when thou divide it my pal, thou get 16 by two two five!

OpenStudy (apoorvk):

looks like \(\large \frac{16}{225}\) - brings back some faded memories aye?

OpenStudy (anonymous):

2 and 6?

OpenStudy (anonymous):

im lost.. again..

OpenStudy (diyadiya):

so what is the probability of picking a black marble = 4/15 we replace it and again pick a black marble (we still have 15 marbles ) so probability of picking black marble again (for the second time) = 4/15 so two probabilities are 4/15 & 4/15 now multiply them \[\frac{4}{15} \times \frac{4}{15}\] i hope you know to do this!!

OpenStudy (diyadiya):

ignore the "so whaat is" in the beginning

OpenStudy (diyadiya):

i was copy-pasting my own posts!^

OpenStudy (anonymous):

i got 16/225

OpenStudy (apoorvk):

you're through!

OpenStudy (diyadiya):

RIGHTTTTTTTT!!!!!!

OpenStudy (anonymous):

thankssssssssssssssssss

OpenStudy (apoorvk):

now start 'rolling' whee-whee

OpenStudy (diyadiya):

You're Welcome ^_^